A contractor needs 96 square feet of brick to construct a rectangular walkway. The length of the walkway is 4 feet less than the width. What is the perimeter of the walkway
Answer by checkley77(12844)

AREA=LW
L=W-4
96=(W-4)W
96=W^2-4W
W^2-4W-96=0
(W-12)(W+8)=0
W-12=0
W=12 FT. FOR THE WIDTH.
L=12-4=8 FT. FOR THE LENGTH.
PROOF:
12*8=96
96=96